OpenStudy (anonymous):

Ok i really need help and fast please!!! So the problem is 3a√ 2a^b^5) can someone SHOW me how to simplify it.

OpenStudy (mertsj):

\[3a \sqrt{2a ^{b ^{5}}}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Oh wait sorry it is 3a√ 2a^7b^5) but i dont know how to eliminate that ^7

OpenStudy (mertsj):

\[3a \sqrt{2a^7b^5}=3a \sqrt{2a^6ab^4b}=3aa^3b^2\sqrt{2ab}=3a^4b^2\sqrt{2ab}\]